# Discrete dark matter with light Dirac neutrinos

### v2

@@ I Introduction
Origin of light neutrino mass and mixing ParticleDataGroup:2020ssz has been one of the longstanding puzzles in particle physics. While neutrinos remain massless within the framework of the standard model (SM) of particle physics, several beyond standard model (BSM) frameworks have been proposed to explain non-zero neutrino mass and mixing. Canonical seesaw frameworks Minkowski:1977sc ; GellMann:1980vs ; Mohapatra:1979ia ; Sawada:1979dis ; Yanagida:1980xy ; Schechter:1980gr ; Mohapatra:1980yp ; Lazarides:1980nt ; Wetterich:1981bx ; Schechter:1981cv ; Foot:1988aq involve inclusion of heavy fields which couple to the SM leptons. These seesaw mechanisms and several of their descendants predict Majorana light neutrino masses which violate lepton number by two units. However, there has been no confirmation about the Majorana nature of light neutrinos with experiments searching for neutrinoless double beta decay (NDBD) continuing to report null results. While null results at NDBD experiments do not necessarily imply Dirac nature of light neutrinos, there has been growing interest in Dirac neutrino models of late. After a few initial attempts in this direction Roncadelli:1983ty ; Roy:1983be ; Babu:1988yq ; Peltoniemi:1992ss , several new proposals to realise sub-eV scale Dirac neutrino masses Chulia:2016ngi ; Aranda:2013gga ; Chen:2015jta ; Ma:2015mjd ; Reig:2016ewy ; Wang:2016lve ; Wang:2017mcy ; Wang:2006jy ; Gabriel:2006ns ; Davidson:2009ha ; Davidson:2010sf ; Bonilla:2016zef ; Farzan:2012sa ; Bonilla:2016diq ; Ma:2016mwh ; Ma:2017kgb ; Borah:2016lrl ; Borah:2016zbd ; Borah:2016hqn ; Borah:2017leo ; CentellesChulia:2017koy ; Bonilla:2017ekt ; Memenga:2013vc ; Borah:2017dmk ; CentellesChulia:2018gwr ; CentellesChulia:2018bkz ; Han:2018zcn ; Borah:2018gjk ; Borah:2018nvu ; Borah:2019bdi ; CentellesChulia:2019xky ; Jana:2019mgj ; Nanda:2019nqy ; Guo:2020qin ; Bernal:2021ezl ; Borah:2022obi ; Li:2022chc ; Dey:2024ctx ; Singh:2024imk have appeared recently. Most of these works, in addition to the inclusion of new fields, also consider new discrete or continuous symmetries in order to forbid unwanted terms in the Lagrangian. In the spirit of Dirac seesaw for light neutrinos, such unwanted terms include the bare mass term of Majorana type for singlet right handed neutrinos $(\nu_{R})$ and direct coupling of $\nu_{R}$ with the SM lepton and Higgs doublets.
@@
Another observed phenomena which can not be explained in the SM is the presence of dark matter (DM) in the Universe. Its presence is supported by astrophysical observations at different scales Zwicky:1933gu ; Rubin:1970zza ; Clowe:2006eq together with cosmological experiments like PLANCK, WMAP predicting around 26.8% of the present Universe to be made up of DM ParticleDataGroup:2020ssz ; Planck:2018vyg . While weakly interacting massive particle (WIMP) has been the most widely studied particle DM framework, null results at direct detection experiments LUX-ZEPLIN:2022qhg have also led to growing interest in alternative scenarios where DM is more feebly coupled to the SM, like the feebly interacting massive particle (FIMP). Recent reviews of WIMP and FIMP can be found in Arcadi:2017kky and Bernal:2017kxu respectively.
@@
Motivated by this, we propose a new flavour symmetric origin of light Dirac neutrino mass and dark matter in this work. We consider the popular non-Abelian discrete flavour symmetry group of $A_{4}$ augmented with an additional discrete $Z_{2}$ and global lepton number $U(1)_{L}$ symmetries. The $A_{4}$ flavour symmetry gets spontaneously broken by flavon fields while generating light Dirac neutrino mass and mixing. For a review of non-Abelian discrete flavour symmetries and their consequences, see Ref. Altarelli:2010gt ; King:2013eh ; Petcov:2017ggy ; Chauhan:2023faf ; Ding:2024ozt and references therein. Unlike popular DM scenarios with additional stabilising symmetries, here we have stable DM by virtue of a remnant $\mathcal{Z}_{2}$ symmetry naturally emerging from $A_{4}$ breaking. Similar scenario for Majorana light neutrino, coined as discrete dark matter was proposed in Hirsch:2010ru and subsequently explored within different contexts Meloni:2010sk ; Boucenna:2011tj ; Adulpravitchai:2011ei ; Eby:2011qa ; Boucenna:2012qb ; Hamada:2014xha ; Lamprea:2016egz ; DeLaVega:2018bkp ; Bonilla:2023pna ; Kumar:2024zfb . We propose a realisation of this discrete dark matter paradigm with light Dirac neutrinos in a minimal $A_{4}$ flavour symmetric setup. We show that neutrino oscillation data can be satisfied after taking tree-level and one-loop contributions to light Dirac neutrino masses. Interestingly, this also leads to the generalisation of the popular Scoto-Seesaw scenario Rojas:2018wym to Dirac neutrinos, and can be termed as Dirac Scoto-Seesaw. While the model predicts either scalar or fermion DM with the desired relic abundance, light Dirac neutrinos lead to observable $\Delta N_{\rm eff}$ at future cosmic microwave background (CMB) experiments like CMB Stage IV (CMB-S4) Abazajian:2019eic , SPT-3G SPT-3G:2014dbx , Simons Observatory SimonsObservatory:2018koc and CMB-HD CMB-HD:2022bsz .
@@ II The Model
In this section, we discuss the relevant field content of the model, the field transformations under the chosen symmetries and the particle spectrum. Table 1 summarises the relevant field content and the corresponding quantum numbers. The SM lepton doublets and right chiral part of Dirac neutrino $(\nu_{R})$ transform as $A_{4}$ singlets while heavy vector-like SM gauge singlet fermions $N_{L,R}$ transform as $A_{4}$ triplets. In addition to the SM Higgs $H$ transforming as $A_{4}$ singlet, additional scalar doublet $\eta$ and scalar singlet $\chi$ both transforming as $A_{4}$ triplets are included. The additional $Z_{2}$ symmetry prevents direct coupling of SM lepton doublet, SM Higgs doublet and $\nu_{R}$. On the other hand, a global unbroken lepton number symmetry $U(1)_{L}$ prevent Majorana mass terms guaranteeing the Dirac nature of light neutrinos. The Yukawa Lagrangian for charged lepton can be written as
@@
with the details of the components $(m_{\nu})_{ij}$ being given in Appendix C. For simplicity, the Yukawa couplings appearing in the neutrino mass matrix are considered to be real. Thus, CP violation is introduced only through the complex parameter $\omega$, appearing here as a consequence of the considered $A_{4}$ discrete flavour symmetry. This mass matrix has rank 3 and predict three non-zero mass eigenvalues, in general. Also, it does not possess any specific symmetric structure and hence can be fitted to neutrino mixing data by appropriate choice of parameters, as we discuss below. Interestingly, in the limit $y_{2}=y_{3}$ and $y^{\prime}_{2}=y^{\prime}_{3}$, the effective mass matrix given in Eq. (11) yields $\mu-\tau$ reflection symmetric mass matrix Harrison:2002et predicting fixed values for $\theta_{23}(=\pi/4)$ and $\delta_{\rm CP}(=\pm\pi/2$), see Appendix C for details. Since type-I seesaw generates only one mass scale with the other being generated by the scotogenic contribution, our framework provides a generalisation of the scoto-seesaw mechanism Rojas:2018wym ; Mandal:2021yph ; Barreiros:2020gxu ; Barreiros:2022aqu ; Ganguly:2022qxj ; Ganguly:2023jml ; Kumar:2023moh to light Dirac neutrinos. Similar to earlier works on Majorana neutrinos, Dirac scoto-seesaw framework can explain the two mass scales in neutrino sector namely, atmospheric and solar mass scales, from tree level (type-I seesaw) and radiative (scotogenic) contributions respectively.
@@
while all other fields are $\mathcal{Z}_{2}$-even. The lightest $\mathcal{Z}_{2}$-odd particle, therefore, is stable and can be a dark matter candidate, the details of which are discussed in section IV.
@@ III Fit to Neutrino data
Figure 2: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\gg y^{\prime}_{i}$ (Case-A).
@@
Figure 3: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\gg y^{\prime}_{i}$ (Case-A).
data Esteban:2020cvm , the allowed ranges for $y_{i},y^{\prime}_{i}$ are given in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes respectively, plotted in Fig. 2. In this benchmark scenario, $y^{\prime}_{i}$ are found to be in the range $6\times 10^{-11}\leq y^{\prime}_{1}\leq 10^{-10}$, $3\times 10^{-11}\leq y^{\prime}_{2,3}\leq 8.5\times 10^{-11}$ and corresponding ranges for $y_{i}$ are found to be $0.03\leq y_{1}\leq 0.05$, $0.052\leq y_{2}\leq 0.1$ and $0.05\leq y_{3}\leq 0.1$ respectively. To be in agreement with neutrino oscillation data, other parameters are restricted within the range $2.8\,\leq(M_{N}/{\rm TeV})\leq 5$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.22\leq(v_{\chi}/{\rm TeV})\leq 1.8$ with the loop functions defined in Eq. (7) having the magnitude $\mathcal{F}_{i}\sim\mathcal{O}(10^{-4})$. With the estimation of the parameters appearing in the neutrino mass matrix, we are now in a position to show the correlation among neutrino oscillation parameters as shown in Fig. 3. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for Case-A and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(0.43-1.6)$ radian, $(0.06-0.09)$ eV and $(8.4-9.7)$ meV respectively.
@@
Case-B: We now consider the benchmark scenario, $y_{i}\ll y^{\prime}_{i}$. Using similar methodology, we can again constrain the Yukawa couplings $y_{i},y^{\prime}_{i}$ imposing the constraints on neutrino masses and mixing angles Esteban:2020cvm .
Figure 4: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\ll y^{\prime}_{i}$ (Case-B).
In Fig. 4 we have plotted the allowed values of the respective Yukawa couplings in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes and $y_{i},y^{\prime}_{i}$ are found to be restricted within the range $1.4\times 10^{-11}\leq y_{1}\leq 2.5\times 10^{-11}$, $2\times 10^{-11}\leq y_{2}\leq 7\times 10^{-11}$, $3.4\times 10^{-11}\leq y_{3}\leq 5.5\times 10^{-11}$
Figure 5: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\ll y^{\prime}_{i}$ (Case-B).
@@
and $0.01\leq y^{\prime}_{1}\leq 0.1$, $0.03\leq y^{\prime}_{2}\leq 0.1$, $0.04\leq y^{\prime}_{3}\leq 0.1$ respectively. In order to satisfy the neutrino oscillation data, other parameters are restricted within the range $6\,\leq(M_{N}/{\rm TeV})\leq 10$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.2\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Corresponding loop functions $\mathcal{F}_{i}$, defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $5\times 10^{-4}\leq\mathcal{F}_{1}\leq 7\times 10^{-4}$, $3\times 10^{-4}\leq\mathcal{F}_{2}\leq 10^{-3}$, $8\times 10^{-4}\leq\mathcal{F}_{3}\leq 1.3\times 10^{-3}$. After estimating the parameters relevant for light neutrino mass matrix, we proceed to find some correlations among neutrino oscillation parameters as shown in Fig. 5. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the scenario $y_{i}\ll y^{\prime}_{i}$ and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range ±$\pm(0-1.6)$ radian, $(0.068-0.11)$ eV and $(8.4-10.5)$ meV respectively.
@@
Case-C: Finally, we now consider the third benchmark scenario where $y_{i}\simeq y^{\prime}_{i}$. In this scenario, one can also satisfy the observed constraints on neutrino masses and mixing
Figure 6: Allowed regions for $y_{i}$, $y^{\prime}_{i}$ for the $3\sigma$ ranges for neutrino oscillation data for $y_{i}\simeq y^{\prime}_{i}$ (Case-C).
where the associated Yukawa couplings are found to be in the range $3\times 10^{-7}\leq y_{1}\leq 4.6\times 10^{-7}$, $6.9\times 10^{-7}\leq y_{2}\leq 7.9\times 10^{-7}$, $7.1\times 10^{-7}\leq y_{3}\leq 8.1\times 10^{-7}$ and $5.6\times 10^{-6}\leq y^{\prime}_{1}\leq 6.8\times 10^{-6}$, $2.9\times 10^{-6}\leq y^{\prime}_{2}\leq 5.5\times 10^{-6}$ and $3.6\times 10^{-6}\leq y^{\prime}_{3}\leq 4.75\times 10^{-6}$ respectively. These allowed ranges are plotted in $y^{\prime}_{i}-y_{i}$ planes as given in Fig. 6.
Figure 7: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\simeq y^{\prime}_{i}$ (Case-C).
@@
To fit the model with the neutrino oscillation data in this scenario, other parameters are restricted within the range $8\,\leq(M_{N}/{\rm TeV})\leq 16$, $35\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.23\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Similarly, loop functions $\mathcal{F}_{i}$ defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $10^{-3}\leq\mathcal{F}_{1}\leq 1.2\times 10^{-3}$, $4\times 10^{-4}\leq\mathcal{F}_{2}\leq 7\times 10^{-4}$, $10^{-4}\leq\mathcal{F}_{3}\leq 1.7\times 10^{-4}$. Similar to the other two cases, here also we show the correlation among neutrino oscillation parameters in Fig. 7. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the current scenario and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(1.2-1.58)$ and $-(1-1.58)$ radian, $(0.082-0.11)$ eV and $(8.8-11)$ meV respectively.

### v3

@@ I Introduction
Origin of light neutrino mass and mixing ParticleDataGroup:2020ssz has been one of the longstanding puzzles in particle physics. While neutrinos remain massless within the framework of the standard model (SM) of particle physics, several beyond the standard model (BSM) frameworks have been proposed to explain non-zero neutrino mass and mixing. Canonical seesaw frameworks Minkowski:1977sc ; GellMann:1980vs ; Mohapatra:1979ia ; Sawada:1979dis ; Yanagida:1980xy ; Schechter:1980gr ; Mohapatra:1980yp ; Lazarides:1980nt ; Wetterich:1981bx ; Schechter:1981cv ; Foot:1988aq involve inclusion of heavy fields which couple to the SM leptons. These seesaw mechanisms and several of their descendants predict Majorana light neutrino masses, which violate the lepton number by two units. However, there has been no confirmation about the Majorana nature of light neutrinos with experiments searching for neutrinoless double beta decay (NDBD) continuing to report null results. While null results at NDBD experiments do not necessarily imply Dirac nature of light neutrinos, there has been growing interest in Dirac neutrino models of late. After a few initial attempts in this direction Roncadelli:1983ty ; Roy:1983be ; Babu:1988yq ; Peltoniemi:1992ss , several new proposals to realise sub-eV scale Dirac neutrino masses Chulia:2016ngi ; Aranda:2013gga ; Chen:2015jta ; Ma:2015mjd ; Reig:2016ewy ; Wang:2016lve ; Wang:2017mcy ; Wang:2006jy ; Gabriel:2006ns ; Davidson:2009ha ; Davidson:2010sf ; Bonilla:2016zef ; Farzan:2012sa ; Bonilla:2016diq ; Ma:2016mwh ; Ma:2017kgb ; Borah:2016lrl ; Borah:2016zbd ; Borah:2016hqn ; Borah:2017leo ; CentellesChulia:2017koy ; Bonilla:2017ekt ; Memenga:2013vc ; Borah:2017dmk ; CentellesChulia:2018gwr ; CentellesChulia:2018bkz ; Han:2018zcn ; Borah:2018gjk ; Borah:2018nvu ; Borah:2019bdi ; CentellesChulia:2019xky ; Jana:2019mgj ; Nanda:2019nqy ; Guo:2020qin ; Bernal:2021ezl ; Borah:2022obi ; Li:2022chc ; Dey:2024ctx ; Singh:2024imk have appeared recently. Most of these works, in addition to the inclusion of new fields, also consider new discrete or continuous symmetries in order to forbid unwanted terms in the Lagrangian. In the spirit of the Dirac seesaw for light neutrinos, such unwanted terms include the bare mass term of Majorana type for singlet right handed neutrinos $(\nu_{R})$ and direct coupling of $\nu_{R}$ with the SM lepton and Higgs doublets.
@@
Another observed phenomenon which can not be explained in the SM is the presence of dark matter (DM) in the Universe. Its presence is supported by astrophysical observations at different scales Zwicky:1933gu ; Rubin:1970zza ; Clowe:2006eq together with cosmological experiments like PLANCK, WMAP predicting around 26.8% of the present Universe to be made up of DM ParticleDataGroup:2020ssz ; Planck:2018vyg . While weakly interacting massive particle (WIMP) has been the most widely studied particle DM framework, null results at direct detection experiments LUX-ZEPLIN:2022qhg have also led to growing interest in alternative scenarios where DM is more feebly coupled to the SM, like the feebly interacting massive particle (FIMP). Recent reviews of WIMP and FIMP can be found in Arcadi:2017kky and Bernal:2017kxu respectively.
@@
Motivated by this, we propose a new flavour symmetric origin of light Dirac neutrino mass and dark matter in this work. We consider the popular non-Abelian discrete flavour symmetry group of $A_{4}$ augmented with an additional discrete $Z_{2}$ and global lepton number $U(1)_{L}$ symmetries. The $A_{4}$ flavour symmetry gets spontaneously broken by flavon fields while generating light Dirac neutrino mass and mixing. For a review of non-Abelian discrete flavour symmetries and their consequences, see Ref. Altarelli:2010gt ; King:2013eh ; Petcov:2017ggy ; Chauhan:2023faf ; Ding:2024ozt and references therein. Unlike popular DM scenarios with additional stabilising symmetries, here we have stable DM by virtue of a remnant $\mathcal{Z}_{2}$ symmetry naturally emerging from $A_{4}$ breaking. Similar scenario for Majorana light neutrino, coined as discrete dark matter was proposed in Hirsch:2010ru and subsequently explored within different contexts Meloni:2010sk ; Boucenna:2011tj ; Adulpravitchai:2011ei ; Eby:2011qa ; Boucenna:2012qb ; Hamada:2014xha ; Lamprea:2016egz ; DeLaVega:2018bkp ; Bonilla:2023pna ; Kumar:2024zfb . We propose a first realisation of this discrete dark matter paradigm with light Dirac neutrinos in a minimal $A_{4}$ flavour symmetric setup. We show that neutrino oscillation data can be satisfied after taking tree-level and one-loop contributions to light Dirac neutrino masses. Interestingly, this also leads to the first generalisation of the popular Scoto-Seesaw scenario Rojas:2018wym to Dirac neutrinos, and can be termed as Dirac Scoto-Seesaw. While the model predicts either scalar or fermion DM with the desired relic abundance, light Dirac neutrinos lead to observable $\Delta N_{\rm eff}$ at future cosmic microwave background (CMB) experiments like CMB Stage IV (CMB-S4) Abazajian:2019eic , SPT-3G SPT-3G:2014dbx , Simons Observatory SimonsObservatory:2018koc and CMB-HD CMB-HD:2022bsz .
@@ II The Model
In this section, we discuss the relevant field content of the model, the field transformations under the chosen symmetries and the particle spectrum. Table 1 summarises the relevant field content and the corresponding quantum numbers. The SM lepton doublets and right chiral part of Dirac neutrino $(\nu_{R})$ transform as $A_{4}$ singlets while heavy vector-like SM gauge singlet fermions $N_{L,R}$ transform as $A_{4}$ triplets. In addition to the SM Higgs $H$ transforming as $A_{4}$ singlet, additional scalar doublet $\eta$ and scalar singlet $\chi$ both transforming as $A_{4}$ triplets are included. The additional $Z_{2}$ symmetry prevents direct coupling of SM lepton doublet, SM Higgs doublet and $\nu_{R}$. On the other hand, a global unbroken lepton number symmetry $U(1)_{L}$ prevent Majorana mass terms from guaranteeing the Dirac nature of light neutrinos. It should be noted that Dirac neutrino seesaw models typically require more symmetries and field content compared to the Majorana neutrino seesaw models due to the requirements of keeping lepton number conserved and the fact that Dirac neutrino has twice the degrees of freedom of its Majorana counterpart. Our choice of symmetry and field content is however not unique. One can replace global lepton number symmetry by discrete symmetries like $Z_{4}$ CentellesChulia:2016rms as well. On the other hand, $Z_{2}$ symmetry ensures seesaw origin of Dirac neutrino mass CentellesChulia:2016rms in such a minimal setup. One can embed such additional discrete symmetries within gauge symmetries as well, but those scenarios will have more particle content. Here we stick to the most minimal extension of the discrete dark matter model Hirsch:2010ru to realise light Dirac neutrinos.
@@
with the details of the components $(m_{\nu})_{ij}$ being given in Appendix C. For simplicity, the Yukawa couplings appearing in the neutrino mass matrix are considered to be real. Thus, CP violation is introduced only through the complex parameter $\omega$, appearing here only in the radiative contribution as a consequence of the considered $A_{4}$ discrete flavour symmetry. This mass matrix has rank 3 and predicts three non-zero mass eigenvalues, in general. Also, it does not possess any specific symmetric structure and hence can be fitted to neutrino mixing data by appropriate choice of parameters, as we discuss below. Interestingly, in the limit $y_{2}=y_{3}$ and $y^{\prime}_{2}=y^{\prime}_{3}$, the effective mass matrix given in Eq. (11) yields $\mu-\tau$ reflection symmetric mass matrix Harrison:2002et predicting fixed values for $\theta_{23}(=\pi/4)$ and $\delta_{\rm CP}(=\pm\pi/2$), see Appendix C for details. Since type-I seesaw generates only one mass scale with the other being generated by the scotogenic contribution, our framework provides a generalisation of the scoto-seesaw mechanism Rojas:2018wym ; Mandal:2021yph ; Barreiros:2020gxu ; Barreiros:2022aqu ; Ganguly:2022qxj ; Ganguly:2023jml ; Kumar:2023moh to light Dirac neutrinos. Similar to earlier works on Majorana neutrinos, the Dirac scoto-seesaw framework can explain the two mass scales in the neutrino sector namely, atmospheric and solar mass scales, from tree level (type-I seesaw) and radiative (scotogenic) contributions, respectively.
@@
while all other fields are $\mathcal{Z}_{2}$-even. Therefore, the lightest $\mathcal{Z}_{2}$-odd particle is stable and can be a dark matter candidate, the details of which are discussed in section IV.
@@ III Fit to Neutrino data
Figure 2: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\delta_{\rm CP}-\sin^{2}\theta_{13}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\gg y^{\prime}_{i}$ (Case-A).
In the subsequent sections, we will explore the viability of these three benchmark scenarios to obtain dark matter relic density, direct search constraints and $\Delta N_{\rm eff}$ elucidating a flavour symmetric origin for these observables. For brevity here we explicitly show the analysis for the normal hierarchy, however such an analysis can also be obtained for inverted hierarchy of the neutrino masses.
@@
Case-B: We now consider the benchmark scenario, $y_{i}\ll y^{\prime}_{i}$. Using a similar methodology, we can again constrain the Yukawa couplings $y_{i},y^{\prime}_{i}$, imposing the constraints on neutrino masses and mixing angles Esteban:2020cvm . The Yukawa couplings $y_{i},y^{\prime}_{i}$ are found to be restricted within the range $1.4\times 10^{-11}\leq y_{1}\leq 2.5\times 10^{-11}$, $2\times 10^{-11}\leq y_{2}\leq 7\times 10^{-11}$, $3.4\times 10^{-11}\leq y_{3}\leq 5.5\times 10^{-11}$
Figure 3: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\delta_{\rm CP}-\sin^{2}\theta_{13}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\ll y^{\prime}_{i}$ (Case-B).
@@
and $0.01\leq y^{\prime}_{1}\leq 0.1$, $0.03\leq y^{\prime}_{2}\leq 0.1$, $0.04\leq y^{\prime}_{3}\leq 0.1$ respectively. In order to satisfy the neutrino oscillation data, other parameters are restricted within the range $6\,\leq(M_{N}/{\rm TeV})\leq 10$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.2\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Corresponding loop functions $\mathcal{F}_{i}$, defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $5\times 10^{-4}\leq\mathcal{F}_{1}\leq 7\times 10^{-4}$, $3\times 10^{-4}\leq\mathcal{F}_{2}\leq 10^{-3}$, $8\times 10^{-4}\leq\mathcal{F}_{3}\leq 1.3\times 10^{-3}$. After estimating the parameters relevant to the light neutrino mass matrix, we proceed to find some correlations among neutrino oscillation parameters as shown in Fig. 3. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\delta_{\rm CP}-\sin^{2}\theta_{13}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the scenario $y_{i}\ll y^{\prime}_{i}$ and predicted ranges for $\delta_{\rm CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $-1.6\leq(\delta_{\rm CP}/{\rm radian})\leq 1.6$, $(0.068-0.11)$ eV and $(8.4-10.5)$ meV respectively.
@@
Case-C: Finally, we now consider the third benchmark scenario where $y_{i}\simeq y^{\prime}_{i}$. In this scenario, one can also satisfy the observed constraints on neutrino masses and mixing where the associated Yukawa couplings are found to be in the range $3\times 10^{-7}\leq y_{1}\leq 4.6\times 10^{-7}$, $6.9\times 10^{-7}\leq y_{2}\leq 7.9\times 10^{-7}$, $7.1\times 10^{-7}\leq y_{3}\leq 8.1\times 10^{-7}$ and $5.6\times 10^{-6}\leq y^{\prime}_{1}\leq 6.8\times 10^{-6}$, $2.9\times 10^{-6}\leq y^{\prime}_{2}\leq 5.5\times 10^{-6}$ and $3.6\times 10^{-6}\leq y^{\prime}_{3}\leq 4.75\times 10^{-6}$ respectively.
Figure 4: Correlation for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\delta_{\rm CP}-\sin^{2}\theta_{13}$, $\sum m_{i}-m_{1}$ and $\sum m_{i}-\sin^{2}\theta_{23}$ for $y_{i}\simeq y^{\prime}_{i}$ (Case-C).
@@
To fit the model with the neutrino oscillation data in this scenario, other parameters are restricted within the range $8\,\leq(M_{N}/{\rm TeV})\leq 16$, $35\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.23\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Similarly, loop functions $\mathcal{F}_{i}$ defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $10^{-3}\leq\mathcal{F}_{1}\leq 1.2\times 10^{-3}$, $4\times 10^{-4}\leq\mathcal{F}_{2}\leq 7\times 10^{-4}$, $10^{-4}\leq\mathcal{F}_{3}\leq 1.7\times 10^{-4}$. Similar to the other two cases, here also we show the correlation among neutrino oscillation parameters in Fig. 4. In this figure, we have plotted the correlations in $\delta_{\rm CP}-\sin^{2}\theta_{23}$ and $\delta_{\rm CP}-\sin^{2}\theta_{13}$ in the upper plane, while $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ in the lower planes and the predicted ranges for $\delta_{\rm CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $1.01.2\leq(\delta_{\rm CP}/{\rm radian})\leq 1.58,-1.58\leq(\delta_{\rm CP}/{\rm radian})\leq-$ 1.0, $\sum m_{i}\in(0.082-0.11)$ eV and $m_{1}\in(8.8-11)$ meV respectively.
It is worth mentioning that the product of the two Yukawa couplings $y_{i}y^{\prime}_{j}$ remain around $\lesssim\mathcal{O}(10^{-12})$ in all the cases mentioned above. This is particularly due to the chosen scale of seesaw $\sim\mathcal{O}(10)$ TeV, singlet flavon VEV around or above the electroweak scale and fine-tuning in the loop function at the level of $\mathcal{O}(10^{-4})$. Unlike in canonical seesaw models where Dirac Yukawa couplings can be large for a very high seesaw scale, scoto-seesaw models offer somewhat less freedom to do so as the same heavy particle masses and Dirac Yukawa couplings take part in tree-level and one-loop contribution to seesaw masses. In the spirit of the scoto-seesaw mechanism both of these contributions are needed to satisfy neutrino data as tree-level seesaw alone gives a neutrino mass matrix of rank-1 only. We check that any fine-tuning of Yukawa couplings can be totally avoided at the expense of fine-tuning the loop function to $\lesssim\mathcal{O}(10^{-14})$. For a given seesaw scale, such fine-tuned loop function can be achieved by choosing tiny scalar quartic couplings involved in $\eta-\chi$ interactions. We, however, do not explore such possibilities. Additionally, the scale of seesaw around a few tens of TeV is chosen keeping in mind the upper limit on thermal dark matter mass which we will discuss in the next section.